Bolts holding a flanged joint in a steam piping are tightened to an initial stress of 400MN/m2. If the bolts are made from steel with a young's modulus of 200GN/m2, determine the relaxed a) stress after 10,000 hours of use. [The creep parameters for this material MN/m2] follows: n 3, A 4.8 x 103per hour per are as b) A series of creep test on an austenitic high temperature steel gave the following results Stress (MNm2) Eo(%) Steady state creep rate 


1
Expert's answer
2021-10-04T04:55:20-0400

"\mathring{\epsilon}=A\sigma^{n}e^{\frac{-Q}{RT}}"

Also "\mathring{\epsilon}=\frac{1}{t}"

Thus, "\frac{1}{t}=A\sigma^{n}e^{\frac{-Q}{RT}}"

n=3

A=4.8*103

Q=136000

R=8.314

T=393

t=10000

"\frac{1}{10000}=4.8*1^{3}\sigma^{3}e^{(\frac{-136000}{8.314*393})}"

"\frac{1}{10000}=4.0223*10^{-15}\sigma^{n3}"

"\sigma^{3}=\frac{1}{10000*4.0223*10^{-15}}=2.4861*10^{10}"

"\sigma=\sqrt[3]{2.4861*10^{10}}"

"=2981.59MPa"